OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1). 1 parallel: y = -3x + 6 perpendicular: y = x + 2 parallel: y = -3x + 10 perpendicular: y = x 3 parallel: y = x perpendicular: y = -3x 4 parallel: y = x + 1 perpendicular: y = 3x - 1

OpenStudy (rmrjr22):

where are the equations?

OpenStudy (anonymous):

o parallel: y = -3x + 6 perpendicular: y = x + parallel: y = -3x + 10 perpendicular: y = x parallel: y = x perpendicular: y = -3x parallel: y = x + 1 perpendicular: y = 3x - 1

OpenStudy (rmrjr22):

parallel usually has same slope perpendicular you have to flip the slope and change the sign

OpenStudy (anonymous):

so it would be #2

OpenStudy (rmrjr22):

seems legit

OpenStudy (anonymous):

aight thanks

OpenStudy (rmrjr22):

is that right?